OpenStudy (anonymous):

find the value of X for which ABCD must be parallelogram 5x-8 , 2x+7

OpenStudy (raden):

diagram ?

OpenStudy (raden):

BC = AD 5x - 8 = 2x + 7 solve for x

OpenStudy (anonymous):

so the answer is x= 6?

OpenStudy (raden):

you just guessing right ?

OpenStudy (anonymous):

no lol?

OpenStudy (raden):

but, that's wrong :P

OpenStudy (anonymous):

whats the answer :(

OpenStudy (raden):

5x - 8 = 2x + 7 first, add by 8 to both sides. what you get

OpenStudy (anonymous):

lol i got 15?

OpenStudy (raden):

ha ? --"

OpenStudy (anonymous):

i dont get ittttt. that's why i need the answer lol

OpenStudy (raden):

5x - 8 = 2x + 7 if you add by 8 to both sides, it means : 5x - 8 + 8 = 2x + 7 + 8 5x = 2x + 15 understand ?

OpenStudy (anonymous):

Ohhhhhh lol. i kinda get it.

OpenStudy (anonymous):

so is the answer 22? or nah?

OpenStudy (raden):

then both sides subtracted by 2x, it means : 5x = 2x + 15 5x - 2x = 2x + 15 - 2x 3x = 15

OpenStudy (anonymous):

5?

OpenStudy (raden):

3 times what be equal 15 ?

OpenStudy (raden):

yes, x = 5

OpenStudy (anonymous):

i need help with another! lol

OpenStudy (anonymous):

is it the same way?

OpenStudy (raden):

i want, but i must going to bed, my eyes just like the lamp 5 watt's :) hint, yes do same idea like before. good luck

OpenStudy (anonymous):

lol ok thanks!